# Conforming Spline-Based Parameterisation Techniques for Plane Graphs

### v1

@@ 2.1 Removing concave corners
4:    while $F_{\text{conc}}^{\varepsilon_{\angle}}(\mathcal{F})\setminus\mathcal{F}_{\text{manual}}\neq\emptyset$ do
5:         $F\leftarrow$ first element of $F_{\text{conc}}^{\varepsilon_{\angle}}(\mathcal{F})\setminus\mathcal{F}_{\text{manual}}$
6:         $C_{\text{found}}\leftarrow\text{False}$
7:         $i\leftarrow 1$
8:         while $i\leq|\mathbb{V}(F)|$ and $C_{\text{found}}$ is False do
9:             $v_{\alpha}\leftarrow$ $i$-th vertex of $\mathbb{V}(F)$
10:             $V_{\text{valid}}\leftarrow\{v_{\beta}\in\mathbb{V}(F)\setminus\{v_{\alpha}\}\,|\,\text{the Hermite splitting curve }C_{\alpha\beta}\text{ is valid}\}$
11:             if $V_{\text{valid}}\neq\emptyset$ then
12:                 $C_{\text{found}}\leftarrow$ True
13:                 $v_{\beta}\leftarrow$ the vertex $v\in V_{\text{valid}}$ that minimizes $\mathcal{Q}^{\mu}(C_{\alpha\beta})$
14:                 $\{F^{+},F^{-}\}\leftarrow$ the two faces that result from splitting $F$ along edge $e_{\alpha\beta}$
15:                 Update $\mathcal{F}\leftarrow\mathcal{F}\setminus\{F\}\cup\{F^{+},F^{-}\}$
16:                 Update $E\leftarrow E\cup\{e_{\alpha\beta}\}$
17:             $i\leftarrow i+1$
18:         if $C_{\text{found}}$ is False then
19:             Update $\mathcal{F}_{\text{manual}}\leftarrow\mathcal{F}_{\text{manual}}\cup\{F\}$              return $\mathcal{F}_{\text{manual}}$
Figure 3: An example showing the splitting of a face with concave corners (blue) by drawing a Hermite curve from a concave corner (red) to another vertex (red or black). Algorithm 1 selects vertex pairs $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F_{i})$, where $v_{\alpha}$ is a concave vertex in a greedy fashion and connects them using a Hermite curve, thus splitting the face in two. The same routine is then applied to the updated plane graph until no more eligible vertex pairs are found. In this example, the routine succeeds in fully automatically removing all concave vertices.
@@ 2.2 Templatisation
    An automated strategy for selecting an appropriate template $T_{i}\in\mathbb{T}$ for an even-sided face $F_{i}\in\mathcal{F}$.
@@
Apart from the obvious filtering step that keeps only the templates $T_{i}\in\mathbb{T}$ with $|\partial E_{i}|=|F_{i}|$, a second filter retains only templates that are compatible with the corner pattern of $F_{i}\in\mathcal{F}$. More precisely, if $|\angle(v)-\pi|$, with $v\in\mathbb{V}(F_{i})$, is small in $F_{i}$, we retain only templates $T_{i}\in\mathbb{T}$ that result in a correspondence wherein the vertex $\phi^{-1}(v)=\hat{v}\in\mathbb{V}(\partial E_{i})$ has valence $\operatorname{val}(\hat{v})\geq 3$ to avoid (nearly) singular corners.
In rare cases (and depending on the template selection strategy), in step C) the ordered set of eligible templates may be empty. This is in particular true for $N$-sided faces with $N>16$. In this case, by default, we fall back on the so-called radial-centric (RC) $N$-leaf template. For $N\geq 6$ the RC $N$-leaf template creates an $N$-sided regular polygon centered in the origin whose vertices are ordered in counterclockwise orientation. Then, it moves a copy $\hat{v}_{i}^{\prime}$ of each boundary node $\hat{v}_{i}$, numbered in counterclockwise orientation, radially inward while connecting them by linear edges. Then, edges are drawn between $\hat{v}_{i}^{\prime}$ and $\hat{v}_{i+1}^{\prime}$ (taking subscripts modulo $N$). This creates a layer of boundary patches. Finally, a node $\hat{v}^{c}$ is placed in the origin and each node $\hat{v}_{i}^{\prime}$, with $i$ odd, is connected to $\hat{v}^{c}$ thus finalising the template quadrangulation. The RC $N$-leaf template with $N=8$ along with its PAG $P(8,12)$ is depicted in Figure 2 of Section 1.6. The procedure for $N=4$ is the same but terminates after the second step (see Figure 4).
Figure 4: The radial-centric four-leaf template.
@@
Figure 5: Figure that visualises one step in the creation of the TSG using the techniques from Section 1.5. The skeleton (center) of the dense polygon corresponding to a face $F_{i}\in\mathcal{F}$ (left) is triangulated and the triangular mesh’s vertices are mapped into the interior of template $T_{i}$ (right). Then, the template’s vertices are mapped back into the skeleton in order to create a covering for it. The vertices that have been added in order to acquire a finer triangulation are highlighted in red.
@@
Having introduced the TSG, we are in the position to introduce the concept of template refinement, a critical aspect of point B) of this section’s three main ingredients. Given an edge $e\in E$, let $F_{e}\subseteq\mathcal{F}$ be the set of faces $F_{i}\in\mathcal{F}$ with |$|\{e,-e\}\cap F_{i}|\neq 0$ and let $F_{\mathcal{T}}:=\{F_{i}\in\mathcal{F}\,|\,T_{i}\in\mathcal{T}\}$ be the subset of templated faces. Clearly, we have $|F_{e}|\in\{1,2\}$. Template refinement enables creating a new vertex by splitting an existing edge $e\in E$ while automatically performing operations on the graph $G=(V,E,\mathcal{F},\mathcal{T})$ such that a suitable template structure is retained in the case $F_{e}\cap F_{\mathcal{T}}\neq\emptyset$. By $G^{\square}=(V^{\square},E^{\square},\mathcal{F}^{\square})$ we denote the canonical TSG of $G$ with quadrangular faces $q^{\square}\in\mathcal{F}^{\square}$. If $|\mathcal{T}|\neq|\mathcal{F}|$, we note that $G^{\square}$ may no longer be connected. Each $q^{\square}\in\mathcal{F}^{\square}$ is associated with a unique $q\in\mathcal{Q}_{i}$ of some $\mathcal{T}\ni T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$. This induces the bijective and exhaustive correspondence
$\Psi^{\square}:\mathcal{F}^{\square}\rightarrow\bigcup_{\mathclap{\begin{subarray}{c}\{\mathcal{Q}_{i}\,|\,T_{i}\in\mathcal{T}\}\end{subarray}}}\mathcal{Q}_{i}.$
Let $e\in E$ satisfy $F_{e}\cap\mathcal{F}_{\mathcal{T}}\neq 0$. Drawing inspiration from classical quadrangulations, template refinement performs classical mesh refinement on the $q^{\square}\in\mathcal{F}^{\square}$ adjacent to the root edge $e\in(E\cap E^{\square})$ while continuing in both directions until boundary cells are encountered. Let $\mathcal{Q}^{\square}_{e}\subseteq\mathcal{F}^{\square}$ and $E_{e}^{\square}\subseteq E^{\square}$ denote the subset of cells and edges, respectively, associated with this operation. For each $q^{\square}\in\mathcal{Q}^{\square}_{e}$, we refine the cell $\Psi^{\square}(q^{\square})\in\mathcal{Q}_{i}$ in the template $T_{i}\in\mathcal{T}$ by halving it in the direction that corresponds to the refinement direction of $q^{\square}\in\mathcal{Q}^{\square}_{e}$. We note that this operation retains the quadrangular nature of the $T_{i}\in\mathcal{T}$ without affecting the associated polygon $\hat{\Omega}_{i}$. We update $G$ by replacing each $e\in(E^{\square}_{e}\cap E)$, where $\iota(e)=(i_{1},i_{2})$, by the two edges $(e^{-},e^{+})$ with $\iota(e^{-})=(i_{1},\alpha)$ and $\iota(e^{+})=(\alpha,i_{2})$ while introducing the new vertex $v_{\alpha}=p$ with $p\in w(e)$. Unless stated otherwise, we choose $p$ as the point that splits the length of the ordered point set $w(e)$ most evenly (as measured by the total length). This creates the associated point sets $(p^{-},p^{+})$ with the assumptions from Section 1, no new corner is created at the shared vertex of $(p^{-},p^{+})$. We finalise the operation by appropriately updating the $F_{i}\in\mathcal{F}$ and the weight function $w(\,\cdot\,)$.
@@
This concept is visualised in Figure 7 which adopts the TSG from Figure 6 (right) and performs template refinement on the edge $e\in(E\cap E^{\square})$ highlighted in blue. The figure furthermore shows the refinement path and the updated TSG and highlights the newly created vertices $v\in V$ of the updated graph $G=(V,E,\mathcal{F},\mathcal{T})$ while furthermore showing the two affected templates before and after refinement. Finally, the patchwise smooth controlmap $\mathbf{r}:{\hat{\Omega}}_{i}\rightarrow{\hat{\Omega}}^{\mathbf{r}}_{i}$ need not be updated because refinement does not change the corner pattern of the updated $F_{i}\in\mathcal{F}$ by assumption.
@@
3:     $\mathcal{T}_{\text{select}}\leftarrow\left\{T\in\mathbb{T}\,\big{|}\,|\partial E|=|F|\right\}$
@@
Type 2. strategies make concrete predictions about the performance of some $\mathbb{T}_{\text{select}}\ni T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$, where $\mathbb{T}_{\text{select}}$ is returned by Algorithm 3, when employed in this paper’s parameterisation method based on harmonic maps. Let $F\in\mathcal{F}$ be a face and let ${\Omega}$ be the associated domain with ${\hat{\Omega}}^{\mathbf{r}}$ the canonical control domain (see Section 1.5). To enable testing a large number of templates, we utilise Floater’s algorithm to approximate a harmonic map between the polygons ${\Omega}_{h}$ and ${\hat{\Omega}}^{\mathbf{r}}_{h}$, both (relatively) coarsely sampled from $\partial{\Omega}$ and $\partial{\hat{\Omega}}^{\mathbf{r}}(F)$, respectively. We sample $N_{\text{sample}}$ points from each edge of $\partial\Omega$ and $\partial{\hat{\Omega}}^{\mathbf{r}}$. The polygon ${\Omega}_{h}$ is triangulated using Gmsh [33] and a surrogate map $\mathbf{x}_{h}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}_{h}\rightarrow{\Omega}_{h}$ is constructed from the correspondence between ${\Omega}_{h}$ and ${\hat{\Omega}}^{\mathbf{r}}_{h}$ using the techniques from Section 1.5.
@@
While more computationally demanding than type 1. strategies, type 2. strategies remain relatively inexpensive because the map $\mathbf{x}_{h}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}_{h}\rightarrow{\Omega}_{h}$ as well as the harmonic map used for the computation of a controlmap $\mathbf{r}(T_{i}):{\hat{\Omega}}(|F|)\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ (see Section 1.6) can be reused.
@@
A more sophisticated strategy adds a further optimisation stage to the controlmap ${\mathbf{r}}:{\hat{\Omega}}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$. Let $T_{i}=(E_{i},V_{i},\mathcal{Q}_{i})$ be an eligible template along with its canonical control template $T^{\mathbf{r}}=(V^{\mathbf{r}},E^{\mathbf{r}},\mathcal{Q}^{\mathbf{r}})$, which follows from replacing $V_{i}\ni\hat{v}\rightarrow{\mathbf{r}}(\hat{v})$ (see Section 1.5). Regarding the controlmap ${\mathbf{r}}:{\hat{\Omega}}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ as a function of the correspondence $V_{i}\ni\hat{v}\rightarrow v^{\mathbf{r}}\in V^{\mathbf{r}}$, we may further optimise the positions of the $v^{\mathbf{r}}\in V^{\mathbf{r}}$ in a way that reduces the value of $C_{T}(T_{i})$. We define the preferred angle in a way analogous to (27):

### v2

@@ 2.1 Removing concave corners
4:    while $\mathcal{F}_{\text{conc}}^{\varepsilon_{\angle}}(\mathcal{F})\setminus\mathcal{F}_{\text{manual}}\neq\emptyset$ do
5:         $F\leftarrow$ first element of $\mathcal{F}_{\text{conc}}^{\varepsilon_{\angle}}(\mathcal{F})\setminus\mathcal{F}_{\text{manual}}$
6:         $E_{\text{valid}}\leftarrow$ the set of $e_{\alpha\beta}$ ending in at least one $v\in\mathcal{V}_{\text{conc}}^{\varepsilon_{\angle}}(F)$ while creating a valid splitting curve $C_{\alpha\beta}$
7:         if $E_{\text{valid}}=\emptyset$ then
8:             Update $\mathcal{F}_{\text{manual}}\leftarrow\mathcal{F}_{\text{manual}}\cup\{F\}$
9:         else
10:             $e\leftarrow e_{\alpha\beta}\in E_{\text{valid}}$ such that $C_{\alpha\beta}$ minimises (16)
11:             Update $\mathcal{F}\leftarrow\mathcal{F}\setminus F\cap\{F^{+},F^{-}\}$, $E\leftarrow E\cup e_{\alpha\beta}$ and $w(\,\cdot\,)$              return $\mathcal{F}_{\text{manual}}$
Figure 5: An example showing the splitting of a face with concave corners (blue) by drawing a Hermite curve from a concave corner (red) to another vertex (red or black). Algorithm 1 selects vertex pairs $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F)$, where $v_{\alpha}$ is a concave vertex in a greedy fashion and connects them using a Hermite curve, thus splitting the face in two. The same routine is then applied to the updated plane graph until no more eligible vertex pairs are found. In this example, the routine succeeds in fully automatically removing all concave vertices.
@@ 2.2 Templatisation
    An automated strategy for selecting an appropriate template $T\in\mathbb{T}$ for an even-sided face $F\in\mathcal{F}$.
@@
Apart from the obvious filtering step that keeps only the templates $T_{i}\in\mathbb{T}$ with $|\partial E_{i}|=|F_{i}|$, a second filter retains only templates that are compatible with the corner pattern of $F_{i}\in\mathcal{F}$. More precisely, if $|\angle(v)-\pi|$, with $v\in\mathbb{V}(F_{i})$, is small in $F_{i}$, we retain only templates $T_{i}\in\mathbb{T}$ that result in a correspondence wherein the vertex $\phi^{-1}(v)=\hat{v}\in\mathbb{V}(\partial E_{i})$ has valence $\operatorname{val}(\hat{v})\geq 3$ to avoid (nearly) singular corners, see Figure 6.
Figure 6: Figure showing a geometry (right) with several vertices (highlighted in red) that create an interior angle close to $\theta=\pi$. We retain only templates (left) that assign to vertices $v\in\mathbb{V}(F_{i})$ with $\angle(v)\approx\pi$ some $\hat{v}\in V_{i}$ with $\operatorname{val}(\hat{v})\geq 3$. The figure additionally shows the convex control domain $\widehat{\Omega}_{i}^{\mathbf{r}}$ that is compatible with the corner pattern of $\Omega_{i}$ (see Section 1.5).
In rare cases (and depending on the template selection strategy), in step C), the ordered set of eligible templates may be empty. This is in particular true for $N$-sided faces with $N>16$. In this case, by default, we fall back on the so-called radial-centric (RC) $N$-leaf template. Figure 7 shows the RC $N$-leaf template for $N\in\{4,6,8,10\}$. For $N\geq 12$ the RC template continues the pattern observed for $N\in\{6,8,10\}$.
Figure 7: Figure showing the radial-centric $N$-leaf template for $N\in\{4,6,8,10\}$.
@@
Figure 8: Figure that visualises one step in the creation of the TSG using the techniques from Section 1.5 and A. The skeleton (center) of the dense polygon corresponding to a face $F_{i}\in\mathcal{F}$ (left) is triangulated and the triangular mesh’s vertices are mapped into the interior of template $T_{i}$ (right). Then, the template’s vertices are mapped back into the skeleton in order to create a covering for it. The vertices that have been added in order to obtain a finer triangulation are highlighted in red.
@@
Having introduced the TSG, we are in the position to introduce the concept of template refinement, a critical aspect of point B) of this section’s three main ingredients. Given an edge $e\in E$, let $F_{e}\subseteq\mathcal{F}$ be the set of faces $F_{i}\in\mathcal{F}$ with |$|\{e,-e\}\cap F_{i}|\neq 0$ and let $\mathcal{F}_{\mathcal{T}}:=\{F_{i}\in\mathcal{F}\,|\,T_{i}\in\mathcal{T}\}$ be the subset of templated faces. Clearly, we have $|F_{e}|\in\{1,2\}$. Template refinement enables creating a new vertex by splitting an existing edge $e\in E$ while automatically performing operations on the graph $G=(V,E,\mathcal{F},\mathcal{T})$ such that a suitable template structure is retained in the case $F_{e}\cap F_{\mathcal{T}}\neq\emptyset$. Let $e_{j}\in E$ be an edge with $|F_{e_{j}}|=1$ that we wish to split. Let $F_{i}\in\mathcal{F}_{\mathcal{T}}$ be the face with ±$\pm e_{j}\in F_{i}$. The edge $e_{j}$, is associated with exactly one other $e_{k}\in F_{i}$ where the pairing follows from $T_{i}\in\mathcal{T}$ as follows: if the quad cell $q\in\mathcal{Q}_{i}$ that contains $\partial E_{i}\ni\hat{e}_{j}=\phi_{i}^{-1}(e_{j})$ is refined in the direction orthogonal to $\hat{e}_{j}$, then $e_{k}=\phi(\hat{e}_{k})$, where $\hat{e}_{k}\in\partial E_{i}$ is the other boundary edge that is split under the refinement (see Figure 10). Continuing the pairing on the (if applicable) next templated face, this leads to a sequence of templated edges $e_{j},e_{k},\ldots,e_{\alpha},e_{\beta}$, with $|F_{e}|=2$ for $e\in\{e_{k},\ldots,e_{\alpha}\}$ while $|F_{e_{\beta}}|=1$. Template refinement now splits all edges in the sequence $e_{j},\ldots,e_{\alpha}$ while the $T_{j}\in\mathcal{T}$ associated with the affected faces $F_{j}\in F_{\mathcal{T}}$ are refined accordingly.
Figure 10: Figure showing the concept of assigning a template $T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$ to a face $F_{i}$ and how the assignment pairs edges $\{e_{j},e_{k}\}\subseteq F_{i}$ under the correspondence $\phi_{i}:\partial E_{i}\rightarrow F_{i}$. The figure additionally shows the refinement path in $T_{i}$ originating from the root edge $\hat{e}_{j}=\phi_{i}^{-1}(e_{j})$ which is contained in the root quad cell $q\in\mathcal{Q}_{i}$.
Let $e\in E$ be an edge from the sequence that is refined. Then, we split $w(e)$ at the point $p\in w(e)$ that minimises the length discrepancy of the resulting point sets $P^{-}$ and $P^{+}$. The point $p\in w(e)$ becomes a new vertex in $V$ and $e$ is split into $e^{-}$ and $e^{+}$ with $\iota(e^{-})=(v_{\alpha},p)$ and $\iota(e^{+})=(p,v_{\beta})$.
With the assumptions from Section 1, no new corner is created at the shared vertex of $(P^{-},P^{+})$. We finalise the operation by appropriately updating $G=(V,E,\mathcal{F},\mathcal{T})$ (including the correspondences $\phi_{i}:\partial E_{i}\rightarrow F_{i}$) and the weight function $w(\,\cdot\,)$.
@@
This concept is visualised in Figure 11 which adopts the TSG from Figure 9 (right) and performs template refinement on the edge $e\in E$ highlighted in blue. The figure furthermore shows the refinement path and the updated TSG and highlights the newly created vertices $v\in V$ of the updated graph $G=(V,E,\mathcal{F},\mathcal{T})$ while furthermore showing the two affected templates before and after refinement. Finally, the patchwise smooth controlmap $\mathbf{r}:{\hat{\Omega}}_{i}\rightarrow{\hat{\Omega}}^{\mathbf{r}}_{i}$ need not be updated because refinement does not change the corner pattern of the updated $F_{i}\in\mathcal{F}$ by assumption.
@@
3:     $\mathcal{T}_{\text{select}}\leftarrow\left\{T_{i}\in\mathbb{T}\,\big{|}\,|\partial E_{i}|=|F|\right\}$
@@
Type 2. strategies make concrete predictions about the performance of some $\mathbb{T}_{\text{select}}\ni T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$, where $\mathbb{T}_{\text{select}}$ is returned by Algorithm 3, when employed in this paper’s parameterisation method based on harmonic maps. Let $F\in\mathcal{F}$ be a face and let ${\Omega}$ be the associated domain with ${\hat{\Omega}}^{\mathbf{r}}$ the canonical control domain (see Section 1.5). To enable testing a large number of templates, we utilise Floater’s algorithm to approximate a harmonic map between the polygons ${\Omega}_{h}$ and ${\hat{\Omega}}^{\mathbf{r}}_{h}$, both (relatively) coarsely sampled from $\partial{\Omega}$ and $\partial{\hat{\Omega}}^{\mathbf{r}}(F)$, respectively. We sample $N_{\text{sample}}$ points from each edge of $\partial\Omega$ and $\partial{\hat{\Omega}}^{\mathbf{r}}$. The polygon ${\Omega}_{h}$ is triangulated using Gmsh [10] and a surrogate (approximately harmonic) map $\mathbf{x}_{h}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}_{h}\rightarrow{\Omega}_{h}$ is constructed from the correspondence between ${\Omega}_{h}$ and ${\hat{\Omega}}^{\mathbf{r}}_{h}$ using the techniques from Section 1.5 and A.
Figure 13: Figure depicting the concept of utilising the surrogate map $\mathbf{x}_{h}^{\mathbf{r}}$ between the surrogate domains $\widehat{\Omega}_{h}^{\mathbf{r}}\approx\widehat{\Omega}^{\mathbf{r}}$ and $\Omega_{h}\approx\Omega$ to optimise the inner vertex positions (green) of the template $T^{r}=(V^{r},E^{r},\mathcal{Q}^{r})$. Here, the inner vertex $v^{r}\in V_{\text{int}}^{r}$ positions are chosen such that the patch interface angles created by the quad layout vertices minimises the deviation from the preferred angle (see equation (25)). While the presented strategy has demonstrated effective results in practice, it is important to acknowledge that alternative quality metrics deserve further exploration.
@@
While more computationally demanding than type 1. strategies, type 2. strategies remain relatively inexpensive because the map $\mathbf{x}_{h}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}_{h}\rightarrow{\Omega}_{h}$ needs to be computed only once and can be reused to gauge the quality of a large number of templates $T_{i}\in\mathbb{T}$. Furthermore, the linear operator $A$ associated with the harmonic map employed in the computation of a controlmap $\mathbf{r}(T_{i}):{\hat{\Omega}}(|F|)\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ (see B) can be reused across multiple faces $F_{i}$ that share the same parametric domain ${\Omega}_{i}$. Here, differing control domains ${\hat{\Omega}}^{\mathbf{r}}_{i}\neq{\hat{\Omega}}^{\mathbf{r}}_{j}$ of faces $F_{i}\neq F_{j}$ with ${\hat{\Omega}}_{i}={\hat{\Omega}}_{j}$ merely change the right-hand-sides in the linear problem $Ax=f$ associated with the control map creation.
@@
A more sophisticated strategy adds a further optimisation stage to the controlmap ${\mathbf{r}}:{\hat{\Omega}}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$. Let $T_{i}=(E_{i},V_{i},\mathcal{Q}_{i})$ be an eligible template along with its canonical control template $T^{\mathbf{r}}=(V^{\mathbf{r}},E^{\mathbf{r}},\mathcal{Q}^{\mathbf{r}})$, which follows from replacing $V_{i}\ni\hat{v}\rightarrow{\mathbf{r}}(\hat{v})$ (see B). Regarding the controlmap ${\mathbf{r}}:{\hat{\Omega}}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ as a function of the correspondence $V_{i}\ni\hat{v}\rightarrow v^{\mathbf{r}}\in V^{\mathbf{r}}$, we may further optimise the positions of the $v^{\mathbf{r}}\in V^{\mathbf{r}}$ in a way that reduces the value of $C_{T}(T_{i})$, see Figure 13. We define the preferred angle in a way analogous to (25):